Compute the three forward starting one year rates given the following annually compounded spot rates: r(0,1) = 1.00%, r(0,2) = 1.18%, r(0,3) = 1.25%, and r(0,4) = 1.3%.
Answer #1


1 year forward rate=1.0118^2/1.01-1=1.360%

2 year forward rate=1.0125^3/1.0118^2-1=1.390%

3 year forward rate=1.013^4/1.0125^3-1=1.450%